Determine the concentrations of K2SO4, K+, and SO42- in a solution prepared by dissolving 2.36 x 10-4 g K2SO4 in 1.50 L of waStuck on how to calculate PPM. Can someone please explain this to me!
Answer #1

Given, [k +] = 1.806 X10-6 M and [sol] 9-03 X107 M To convert them from M to ppm first what se ppm, we known bbm → ports perto do this we known the molar mes of Kt Molar mess=kt 39.09 g/mol i moles | 2 of K+ 1.806x10-6 mof x 39.099|reot 70 59 x 10-6moles 12 of kt 70.59 X 10–3 mille 70.59x100 ngle pole | 4 0 4t because ppm and 10-6 ingle = 10-3 10-6 103 ppm = mg/l 3 kt 70.